OpenStudy (anonymous):

find n such that C(n,2)=91

OpenStudy (anonymous):

No such n exists.

OpenStudy (anonymous):

n = 14

OpenStudy (anonymous):

Whoops, haha you're right.

OpenStudy (anonymous):

That's because C(n, 2) = n(n-1)/2. So you want to solve n(n-1)/2 = 91. But then that’s n(n-1) = 182. But factors of 182 are just 2*7*13. So we can see 182 = 13*14, so n = 14.

OpenStudy (anonymous):

good.......

OpenStudy (anonymous):

but something u know not by calculation, by your own experience.......